# Conceptualization, Investigation, Data curation, Formal analysis, Writing - Original draft preparation

### v1

@@ 2.2 Data Collection
We collected the data for our study from GitHub, which is currently the largest platform for hosting open-source software projects. In addition, every project on GitHub has an “Issues” page, which provides a platform for developers and users of the project to raise issues encountered while developing and using the software, and to discuss the identification, tracing, and resolution of these issues with other developers and users.
@@
Following the selection criteria, the first author ultimately selected 15 LLM open-source projects (see Table 1) that met our criteria. To explore the causes of problems and solutions as comprehensively as possible, the first author employed web scrapers to collect all the closed issues from these project repositories. In the end, we gathered a total of 14,588 closed issues. Each closed issue was recorded in an MS Excel file. We named each MS Excel file using the convention: ‘Project Name’ + ‘_’ + ‘Issue Number’.
@@ 2.3 Data Labelling and Sampling
After collecting closed issues from the 15 LLM open-source projects, we conducted data labelling to identify issues suitable for our study. The closed issues we expected should contain specific information related to the issues that practitioners encountered while developing or using LLM open-source software. Then, on the basis of the labelling results (as detailed in Section 2.3.1 and Section 2.3.2), we selected a representative data sample (as detailed in Section 2.3.3) to construct the dataset for data extraction.
@@ 2.3.1 Pilot Data Labelling
To minimize potential personal biases, the first and third authors independently conducted a pilot data labelling. We randomly selected 50 closed issues from all closed issues, and then the first and third authors independently labelled them. The two authors labelled the data that contain information related to issues arising during the development or use of LLM open-source software. The inter-rater consistency on data labelling results between the two authors was measured using the Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.838, which indicates a reasonable level of data labelling consistency between the two authors. For any disagreements or conflicts in the pilot data labelling results, the two authors discussed together with the second author to reach a consensus. Finally, wo got 32 issues out of 50 issues are considered relevant. The results of pilot data labelling were compiled and recorded in MS Excel files (Cai et al., 2024).
@@ 2.3.2 Formal Data Labelling
The first author then conducted the formal data labelling. In this process, we excluded the issues not related to our study. The projects we selected utilized LLMs to achieve the main functions of the software. However, some issues raised by practitioners may not be related to the problems with the projects. For example, some practitioners only raised an issue to chat with others or to learn how the features of GitHub, like a practitioner simply submitted an issue titled “hi” as a greeting, without any content related to the software. We only kept those issues related to the problems that practitioners encountered during the development and use of LLM open-source software. Ultimately, the first author collected a total of 14,476 closed issues. The results of formal data labelling were compiled and recorded in MS Excel files (Cai et al., 2024).
@@ 2.3.3 Data Sampling
We found that the dataset comprising a total of 14,476 closed issues was too large for manual data extraction. Therefore, we decided to select a suitable number of closed issues. We randomly selected 994 closed issues to form a representative set as our dataset for data extraction with a 95% confidence level and a 3% margin of error (Israel, 1992).
@@ 2.4 Data Extraction
To answer the three RQs presented in Section 2.1, we defined a set of data items for data extraction, as shown in Table 2. The data items D1-D3 are respectively used to extract information about issues, underlying causes, and solutions that address these issues, corresponding to RQ1-RQ3. These three data items can be extracted from any part of the content of the collected closed issues, including the title, description, and discussion of the issue.
@@ 2.4.1 Pilot Data Extraction
We randomly selected 50 closed issues from our dataset. Next, the first author independently conducted the pilot data extraction on the 50 issues. The results of pilot data extraction indicated that all three data items (see Table 2) could be extracted from our dataset. Based on the pilot data extraction results, we reached a consensus and defined the following rules for formal data extraction: (1) In principle, only one issue can be extracted from each closed issue. If there are multiple issues that can be extracted, we record the first issue discussed. (2) If there are multiple causes mentioned in a closed issue, we record all the causes identified by the reporter and the development team. (3) If there are multiple solutions mentioned in a closed issue, we record all the solutions identified by the reporter and the development team.
@@ 2.4.2 Formal Data Extraction
Not all closed issues had a corresponding cause and solution which could be extracted. For some closed issues, the discussion content of the issues was too vague to discern specific information (i.e., cause or solution) that could be extracted. Moreover, if an issue of a project is not discussed or updated for a long period, it will be automatically closed and turn into a closed issue, making it impossible to extract valid cause and solution. An issue may also be temporarily closed due to its low priority for resolution. Similarly, an issue might be too complex to identify its causes and solutions.
@@ 2.5 Data Analysis
The steps in our data analysis are as follows: (1) The first author meticulously read through the text of each closed issue, then assigned descriptive codes that succinctly summarize the core themes according to the three RQs. For example, the issue of langchain #2174 was coded as “Model Test Failure”, which indicates that an error occurred while testing the model leading to failure. (2) The first author compared the commonalities and differences between codes. Following the principles of a bottom-up approach, the first author categorized identical or similar codes, and used consistent descriptive language to describe the higher-level types and categories.For example, the code from langchain #2174, together with other similar codes, constitutes the Model Test Issue type, which further belongs to Model Issue category. (3) Once uncertainties about the code descriptions arose, the first author engaged in discussions with the second author to reach a consensus. Due to the nature of Constant Comparison, the initial results underwent several iterations and modifications before reaching the final results. (4) The second author reviewed and verified the initial analysis results. In case of any doubts or disagreements, the first and the second authors engaged in discussions to reach a final consensus and ultimately resolved the conflict.
@@ 4 Implications
4.1 Implications for Practitioners
Implication 1. Practitioners in LLM open-source software development should possess a basic set of knowledge of LLMs.
According to our study results, the following knowledge is required to develop and use LLM open-source software: (1) Practitioners need knowledge of configuring the development and runtime environments for LLM open-source software. According to our study results, Configuration and Connection Problem ranks the second among the causes of the issues, which indicates that practitioners need solid knowledge about environment configuration of LLM open-source software. For example, a contributor received a message that “Package ‘tokenizers’ not found in index and Package ‘taggers’ not found in index”, which indicates key packages are missing from the dependency environment when configuring the development environment for LLM open-source software (langchain #8419). (2) Practitioners need to understand basic text processing techniques to handle input and output texts. LLMs are a kind of PLMs, and their input and output are text. Therefore, when developing LLM open-source software, it is necessary to have a basic understanding of text processing techniques like tokenization, stemming, and lemmatization, stop word removal. For example, a contributor reported that “–no-stream is very slow, because it ignores the stop words filled into stopping_criteria”, which indicates that operational efficiency of LLM open-source software can be enhanced after the removal of stop words (text-generation-webui #805). (3) Practitioners need to have a basic understanding of deep learning frameworks that LLMs (e.g., GPT4) are based on. Acquiring knowledge of deep learning frameworks can facilitate the fine-tuning and inference functionalities of LLMs. For example, a contributor received a message that “RuntimeError: PyTorch is not linked with support for mps devices1”, which indicates that practitioners need a basic understanding of PyTorch to deploy the software (FastChat #854).
Implication 2. Practitioners should use more efficient parameter fine-tuning methods and design more robust parameter validation mechanisms.
According to Table 4, it can be observed that 22 cases of Model Issue are caused by Parameter Problem, 24 cases of Parameter Issue are induced by Model Problem, which reflects that Model Issues and Parameter Issues in LLM open-source software are interrelated, highlighting the critical importance of parameter settings, fine-tuning, and validation for the stability and performance of models. Therefore, practitioners need to introduce more effective parameter tuning methods for LLMs to avoid potential negative impacts on the performance of models. Zhou et al. (2024) suggested employing efficient parameter-efficient fine-tuning (PEFT) methods such as LoRA, IA3, Adapter, and Prefix-Tuning for fine-tuning multimodal models, which may even outperform full fine-tuning (FFT). Additionally, practitioners (e.g., developers of LLMs, OpenAI) should design more robust parameter validation mechanisms to minimize Model Issues arising from incorrect parameter values and types.
@@ 4.1 Implications for Practitioners
Implication 3. Users of LLM open-source software can optimize input prompts to obtain more satisfactory answers from LLM open-source software.
@@
Answer Issue ranks the fourth among all 15 categories of issues, and within Answer Issue, the dominant type is Poor-quality Answer (42.7%), which indicates a level of dissatisfaction among users with the responses provided by LLM open-source software. However, some techniques can be employed to improve the input prompts in order to obtain better responses. For example, users can utilize predefined templates to input prompts for obtaining more refined responses. A user suggested that “you can provide context like below as in the Alpaca data”, which indicates that users can mimic the format of training data by segmenting input prompts into multiple parts such as instructions, input, and response, and write the context according to the syntax and standards of the training data to achieve better responses (dolly #55). Furthermore, users can also divide input prompts that exceed the length limit into multiple segments and enter these segments separately, filter out sensitive vocabulary, or provide more specific scenarios to obtain better responses.
@@
Implication 4. Developers need to be cautious when selecting models and components when developing LLM open-source software.
@@
4.2 Implications for Researchers
Implication 5. Researchers may consider to provide automated parameter adjustment tools and methods.
@@ 4.2 Implications for Researchers
Implication 6. Researchers may consider to provide methods to reproduce the problems occurred in LLM open-source software.
@@
Implication 7. Researchers may consider establishing standardized benchmarks to evaluate the impact of different configurations (e.g., model, dataset, hardware, and software environment) on the performance of LLMs.
@@
Implication 8. Researchers may consider exploring how to better coordinate the collaboration between models and the interoperation between components to reduce incompatibilities within the system.
@@ 5 Threats to Validity
Construct validity: Due to the manual operation of data collection, labelling, sampling, and extraction in our study, there is a risk of personal bias on the data labelling and extraction results. To mitigate this risk and enhance the objectivity and reliability of our study results, we conducted pilot studies before each formal step to assess the effectiveness of methods and criteria among different researchers, which increased the construct validity of the study. Following each pilot and formal study, the first author engaged in discussions with the second and third authors to reach consensus on the data analysis results aligned with the RQs.
@@
Reliability: To mitigate potential uncertainties arising from the research methodologies employed, we implemented several measures to enhance the reliability of our study. As mentioned in Section 2.3.1, we conducted pilot studies to assess consistency on data labelling results between the first and third authors before the formal studies. In the pilot data labelling phase, we computed a Cohen’s Kappa coefficient of 0.838, which indicated good consistency between authors. Throughout the process of data labelling, extraction, and analysis, thorough discussions and resolutions were conducted among the authors to address any internal inconsistencies, ensuring the consistency and accuracy of the results. Furthermore, we have provided the dataset of our study (Cai et al., 2024) for other researchers to replicate this study and validate our findings.

### v2

@@ 2.2 Data Collection
We collected the data for our study from GitHub, which is currently the largest platform for hosting open-source software projects. In addition, every project on GitHub has an “Issues” page, which provides a platform for developers and users of the project to raise issues encountered while developing and using the software, and to discuss the identification, tracing, and resolution of these issues with other developers and users. In order to thoroughly explore the causes of issues and their potential solutions, we decided to collect all “closed” issues from the selected LLM open-source project repositories.
@@
Following the selection criteria, the first author ultimately selected 15 LLM open-source projects (see Table 1) that met our criteria. The domains of the 15 selected LLM open-source projects are categorized into five categories (see Table 2). Then, the first author employed web scrapers to collect all the closed issues from these project repositories. In the end, we gathered a total of 14,588 closed issues. Each closed issue was recorded in an MS Excel file. We named each MS Excel file using the convention: ‘Project Name’ + ‘_’ + ‘Issue Number’.
Table 2: Domain categorization of the selected LLM open-source projects
Domain #Definition #Projects
LLM Development Framework Open-source projects that provide frameworks for developing software by leveraging LLMs langchain, MetaGPT, quivr, autogen, ChatDev, semantic-kernel, SuperAGI, haystack, dolly
LLM Running Tool Open-source projects for running LLMs in specific environments gpt4all, FastChat, ollama
LLM Data Analysis Tool Open-source projects for data analysis using LLMs pandas-ai
LLM Training Tool Open-source projects that support pre-training across tasks, languages, and modalities unilm
LLM Frontend Tool Open-source projects that develop user interface for LLMs text-generation-webui
@@ 2.3 Data Labelling and Sampling
After collecting closed issues from the 15 LLM open-source projects, we found that some issues raised by practitioners may not be related to the problems with the projects. For example, some practitioners only raised an issue to chat with others or to learn the features of GitHub, like a practitioner simply submitted an issue titled “hi” as a greeting, without any content related to the project. In addition, not all issues have corresponding category labels (e.g., pandas-ai issue #85). Relying solely on existing labels could lead to the omission of critical information. Therefore, we conducted data labelling to identify issues suitable for our study. The closed issues we expected should contain specific information related to the issues that practitioners encountered while developing or using LLM open-source software. For example, a user mentioned in text-generation-webui #2926 that “the model is loaded into memory without any errors, but crashes on generation of text”, which explicitly indicates an issue present in the LLM open-source project text-generation-webui. Then, on the basis of the pilot and formal labelling results (as detailed in Section 2.3.1 and Section 2.3.2, respectively), we selected a representative data sample (as detailed in Section 2.3.3) to construct the dataset for data extraction.
@@ 2.3.1 Pilot Data Labelling
To minimize potential personal biases, the first and third authors independently conducted a pilot data labelling. We randomly selected 50 closed issues from all closed issues, and then the first and third authors independently labelled them. The two authors labelled the data that contain information related to issues arising during the development or use of LLM open-source software. The inter-rater consistency on data labelling results between the two authors was measured using the Cohen’s Kappa coefficient (Cohen, 1960), yielding a value of 0.838, which indicates a very high level of data labelling consistency between the two authors. For any disagreements or conflicts in the pilot data labelling results, the two authors discussed together with the second author to reach a consensus. Finally, wo got 32 issues out of 50 issues are considered relevant. The results of pilot data labelling were compiled and recorded in MS Excel files (Cai et al., 2024).
@@ 2.3.2 Formal Data Labelling
The first author then conducted the formal data labelling. In this process, we excluded the issues not related to our study. The projects we selected utilized LLMs to achieve the main functions of the software. We only kept those issues related to the problems that practitioners encountered during the development and use of LLM open-source software. Ultimately, the first author collected a total of 14,476 closed issues. The results of formal data labelling were compiled and recorded in MS Excel files (Cai et al., 2024).
@@ 2.3.3 Data Sampling
We found that the dataset comprising a total of 14,476 closed issues was too large for manual data extraction. Therefore, we decided to select a suitable number of closed issues. We randomly selected 994 closed issues to form a representative set as our dataset for data extraction with a 95% confidence level and a 3% margin of error (Israel, 1992). The source projects of the closed issues in the final data sample are presented in Table 3. In addition, the distribution of the duration from issue opening to closing for the closed issues in the final data sample is shown in Fig. 2.
Figure 2: The distribution of the duration of the closed issues in the final data sample
Table 3: The source projects of the closed issues in the final data sample
Project Name #Selected Issues #Contributors #Lines of Code
langchain 226 1991 1,468.7k
gpt4all 50 79 238.1k
MetaGPT 9 45 133.1k
FastChat 63 206 121.7k
text-generation-webui 178 264 162.9k
ollama 49 89 276.4k
quivr 52 95 61.4k
autogen 19 359 1,037.4k
ChatDev 11 40 1,372k
unilm 50 54 2,047.9k
semantic-kernel 75 195 726.4k
SuperAGI 22 61 62.1k
haystack 162 232 140.3k
dolly 9 13 1.6k
pandas-ai 19 59 84.0k
Total 994
@@ 2.4 Data Extraction
To answer the three RQs presented in Section 2.1, we defined a set of data items for data extraction, as shown in Table 4. The data items D1-D3 are respectively used to extract information about issues, underlying causes, and solutions that address these issues, corresponding to RQ1-RQ3. These three data items can be extracted from any part of the content of the collected closed issues, including the title, description, and discussion of the issue. To ensure the accuracy of the data extraction results and the validity of subsequent data analysis, we decided to conduct the data extraction manually. The other reason is that the data sample size was not very large, and manual data extraction was feasible. Therefore, the data extraction was conducted manually.
@@ 2.4.1 Pilot Data Extraction
We randomly selected 50 closed issues from our dataset. Next, the first author independently conducted the pilot data extraction on the 50 issues. The results of pilot data extraction indicated that all three data items (see Table 4) could be extracted from our dataset. Based on the pilot data extraction results, we reached a consensus and defined the following rules for formal data extraction: (1) In principle, only one issue can be extracted from each closed issue. If there are multiple issues that can be extracted, we record the first issue discussed. (2) If there are multiple causes mentioned in a closed issue, we record all the causes of the issue we extract in the former step, which are identified by the reporter and the development team. (3) If there are multiple solutions mentioned in a closed issue, we record all the solutions identified by the reporter and the development team that can address the issue we extract in the former step.
@@ 2.4.2 Formal Data Extraction
In the formal data extraction phase, we totally got 11 closed issues that discussed multiple issues in one closed issue in the final data sample. However, according to our criteria, we only kept the first issue discussed and its corresponding causes and solutions.
Besides, not all closed issues had a corresponding cause and solution which could be extracted. For some closed issues, the discussion content of the issues was too vague to discern specific information (i.e., cause or solution) that could be extracted. Moreover, if an issue of a project is not discussed or updated for a long period, it will be automatically closed and turn into a closed issue, making it impossible to extract valid cause and solution. An issue may also be temporarily closed due to its low priority for resolution. Similarly, an issue might be too complex to identify its causes and solutions.
@@ 2.5 Data Analysis
The steps in our data analysis are as follows: (1) The first author meticulously read through the text of each closed issue, then assigned descriptive codes that succinctly summarize the core themes according to the three RQs. For example, the issue of langchain #2174 was coded as “Model Test Failure”, which indicates that an error occurred while testing the model leading to failure. (2) The first author compared the commonalities and differences between codes. Following the principles of a bottom-up approach, the first author categorized identical or similar codes, and used consistent descriptive language to describe the higher-level types and categories. For example, the code from langchain #2174, together with other similar codes, constitutes the Model Test Issue type, which further belongs to Model Issue category. (3) Once uncertainties about the code descriptions arose, the first author engaged in discussions with the second author to reach a consensus. Due to the nature of Constant Comparison, the initial results underwent several iterations and modifications before reaching the final results. (4) The second author reviewed and verified the initial analysis results. In case of any doubts or disagreements, the first and the second authors engaged in discussions to reach a final consensus and ultimately resolved the conflict. After resolving 31 conflicts in the categories of issues, 22 conflicts in the categories of causes, and 27 conflicts in the categories of solutions, we obtained the final results of the data analysis.
@@ 4 Implications
4.1 Implications for Users of LLM Open-Source Software
@@
Implication 1. Users of LLM open-source software can optimize input prompts to obtain more satisfactory answers from LLM open-source software.
@@
Answer Issue ranks the fourth among all 15 categories of issues, and within Answer Issue, the dominant type is Poor-quality Answer (42.7%), which indicates a level of dissatisfaction among users with the responses provided by LLM open-source software. However, some techniques can be employed to improve the input prompts in order to obtain better responses. For example, users can utilize predefined templates to input prompts for obtaining more refined responses. A user suggested that “you can provide context like below as in the Alpaca data”, which indicates that users can mimic the format of training data by segmenting input prompts into multiple parts such as instructions, input, and response, and write the context according to the syntax and standards of the training data to achieve better responses (dolly #55). Furthermore, users can also divide input prompts that exceed the length limit into multiple segments and enter these segments separately, filter out sensitive vocabulary, or provide more specific scenarios to obtain better responses. For example, the langchain project provides several built-in text splitting tools, such as CharacterTextSplitter, which can divide a document into smaller chunks for input into LLMs, and allows maintaining continuity between chunks by setting “chunk_overlap”.
Implication 2. Users can consider using LLM Development Frameworks as assistants for their software development, while remaining mindful of the potential problems and risks associated with LLMs.
Over half of the LLM open-source projects investigated in this study are frameworks for developing software by leveraging LLMs (see Table 2), which could lower the barrier for users to develop software systems with the assistance of LLMs. For example, the langchain project provides an open-source software framework for developing applications powered by LLMs, and it supports both open-source and commercial LLMs (such as GPT from OpenAI, Llama from Meta), allowing users of the langchain project to select the LLMs that fit their needs in software development. However, according to the results of our study, there are plenty of issues associated with generating code using LLMs. 46.3% of Model Issues, 71.7% of Component Issues, 65.2% of Parameter Issues, 53.7% of Answer Issues, and 63.4% of Performance Issues originate from LLM Development Frameworks, which indicates that using LLM open-source software for development might lead to many potential problems and risks. For example, Fu et al. (2025) examined the security risks of Copilot-generated code, and found that that 29.5% of Python and 24.2% of JavaScript code snippets contain security weaknesses. Users of LLM Development Frameworks need to be aware of such issues associated with LLMs and take appropriate measures to prevent insecure code merging into the code base.
4.2 Implications for Developers of LLM Open-Source Software
Implication 3. Developers of LLM open-source software should possess a basic set of knowledge of LLMs.
According to our study results, the following knowledge is required to develop and use LLM open-source software: (1) Developers need knowledge of configuring the development and runtime environments for LLM open-source software. According to our study results, Configuration and Connection Problem ranks the second among the causes of the issues, which indicates that practitioners need solid knowledge about environment configuration of LLM open-source software. For example, a contributor received a message that “Package ‘tokenizers’ not found in index and Package ‘taggers’ not found in index”, which indicates key packages are missing from the dependency environment when configuring the development environment for LLM open-source software (langchain #8419). (2) Developers need to understand basic text processing techniques to handle input and output texts. LLMs are a kind of PLMs, and their input and output are text. Therefore, when developing LLM open-source software, it is necessary to have a basic understanding of text processing techniques like tokenization, stemming, and lemmatization, stop word removal. For example, a contributor reported that “–no-stream is very slow, because it ignores the stop words filled into stopping_criteria”, which indicates that operational efficiency of LLM open-source software can be enhanced after the removal of stop words (text-generation-webui #805). (3) Developers need to have a basic understanding of deep learning frameworks that LLMs (e.g., GPT4) are based on. Acquiring knowledge of deep learning frameworks can facilitate the fine-tuning and inference functionalities of LLMs. For example, a contributor received a message that “RuntimeError: PyTorch is not linked with support for mps devices1”, which indicates that practitioners need a basic understanding of PyTorch to deploy the software (FastChat #854).
Implication 4. Developers should use more efficient parameter fine-tuning methods and design more robust parameter validation mechanisms.
According to Table 6, it can be observed that 22 cases of Model Issue are caused by Parameter Problem, 24 cases of Parameter Issue are induced by Model Problem, which reflects that Model Issues and Parameter Issues in LLM open-source software are interrelated, highlighting the critical importance of parameter settings, fine-tuning, and validation for the stability and performance of models. Therefore, developers need to introduce more effective parameter tuning methods for LLMs to avoid potential negative impacts on the performance of models. Zhou et al. (2024) suggested employing efficient parameter-efficient fine-tuning (PEFT) methods such as LoRA, IA3, Adapter, and Prefix-Tuning for fine-tuning multimodal models, which may even outperform full fine-tuning (FFT). Additionally, developers (e.g., developers of LLMs, OpenAI) should design more robust parameter validation mechanisms to minimize Model Issues arising from incorrect parameter values and types. For example, Frantar and Alistarh (2023) proposed a method called SparseGPT, which significantly compresses LLMs by pruning a vast number of parameters. The method uses Hessian information to determine which weights should be pruned and which should be retained. During the pruning process, the pruning mask and weight updates are dynamically adjusted based on the error at each layer and Hessian information to validate whether the pruned parameters negatively affect the accuracy of the LLMs.
@@ 4.2 Implications for Developers of LLM Open-Source Software
Implication 5. Developers need to be cautious when selecting models and components when developing LLM open-source software.
@@
4.3 Implications for Researchers of LLM Open-Source Software
Implication 6. Researchers may consider to provide automated parameter adjustment tools and methods.
@@ 4.3 Implications for Researchers of LLM Open-Source Software
Implication 7. Researchers may consider to provide methods to reproduce the problems occurred in LLM open-source software.
@@
Implication 8. Researchers may consider establishing standardized benchmarks to evaluate the impact of different configurations (e.g., model, dataset, hardware, and software environment) on the performance of LLMs.
@@
Implication 9. Researchers may consider exploring how to better coordinate the collaboration between models and the interoperation between components to reduce incompatibilities within the system.
@@ 5 Threats to Validity
Construct validity: Due to the manual operation of data collection, labelling, sampling, and extraction in our study, there is a risk of personal bias on the data labelling and extraction results. To mitigate this risk and enhance the objectivity and reliability of our study results, we conducted pilot studies before each formal step to assess the effectiveness of methods and criteria among different researchers, which increased the construct validity of the study. Following each pilot and formal study, the first author engaged in discussions with the second and third authors to reach consensus on the data analysis results aligned with the RQs. Since the open coding procedure was carried out by the first author, we acknowledge the potential impact of overconfidence of the first author on the validity of the data extraction results. However, the first author had five years of experiences in software development two years of experience in using LLMs, which partially mitigates the potential threat to the validity of the data extraction results. Besides, we consider adopting automated or semi-automated tools as a future measure to mitigate potential biases in data selection.
@@
Reliability: To mitigate potential uncertainties arising from the research methodologies employed, we implemented several measures to enhance the reliability of our study. As mentioned in Section 2.3.1, we conducted pilot studies to assess consistency on data labelling results between the first and third authors before the formal studies. In the pilot data labelling phase, we computed a Cohen’s Kappa coefficient value of 0.838, which indicated good consistency between authors. Throughout the process of data labelling, extraction, and analysis, thorough discussions and resolutions were conducted among the authors to address any internal inconsistencies, ensuring the consistency and accuracy of the results. Furthermore, we have provided the dataset of our study (Cai et al., 2024) for other researchers to replicate this study and validate our findings.